# Quality-Aware Image-Text Alignment for Real-World Image Quality Assessment

### v2

@@ 4.2 Ablation Studies
To ensure fairness, we conduct ablation studies and tune hyperparameters using the LIVE [33] and TID2013 [27] synthetic datasets, reserving the authentic datasets solely for testing, akin to real-world scenarios. We find that the findings of the ablation studies also hold for the authentic datasets (see the supplementary material for more details).
LIVE TID2013
$\mathcal{L}_{cons}$ $\mathcal{L}_{pos}$ $\mathcal{L}_{neg}$ SRCC PLCC SRCC PLCC
0.602 0.619 0.507 0.617
0.655 0.631 0.519 0.601
0.875 0.856 0.614 0.676
0.674 0.652 0.525 0.618
0.884 0.864 0.629 0.698
0.865 0.863 0.609 0.660
0.893 0.885 0.632 0.705
LIVE TID2013
Ablation SRCC PLCC SRCC PLCC
D = 2 0.844 0.819 0.618 0.695
L = 3 0.816 0.786 0.587 0.679
quality-based 0.831 0.818 0.552 0.681
w/ pos. emb. 0.886 0.872 0.634 0.699
QualiCLIP 0.893 0.885 0.632 0.705
@@
Table 3: Ablation studies on the loss terms of Eq. 4 (top) and on the training strategy (bottom). Best and second-best scores are highlighted in bold and underlined, respectively.
Loss terms We study the importance of each loss term in Eq. 4 and report the results in the top part of Tab. 3. First, we notice that using only $\mathcal{L}_{cons}$ leads to a significant performance decrease, as it does not exploit the information provided by the intrinsic ranking of the increasingly degraded crops. Nevertheless, $\mathcal{L}_{cons}$ consistently yields a positive impact when combined with any of the other loss terms. Then, we observe that despite being analogous to $\mathcal{L}_{pos}$, $\mathcal{L}_{neg}$ seems to be more crucial for the training process. Given that $\mathcal{L}_{neg}$ involves the alignment between the images and the negative prompt, this outcome suggests that such prompt holds more significance in the quality score computation as in Eq. 5. We discuss this hypothesis in the supplementary material. Nevertheless, Tab. 3 shows that combining the three loss terms achieves the best results, proving that they are all crucial for training CLIP to generate accurate quality-aware image representations.
@@
Training strategy We evaluate the performance achieved by modified versions of our approach: 1) $D\!=\!2$: we apply two sequential degradations to each crop in Sec. 3.2 instead of just one; 2) $L\!=\!3$: we consider only three levels of degradation in Secs. 3.2 and 3.3 instead of five; 3) we compute the ranking loss using the predicted quality scores – obtained with Eq. 5 – associated to each degraded crop instead of its similarity to the antonym prompts; 4) w/ pos. emb.: we relax CLIP’s requirement of fixed-size inputs by interpolating the positional embedding instead of removing it. The bottom part of Tab. 3 shows the results. First, we note that employing more than one distortion leads to worse performance. This stems from the synthetic degradation becoming too severe independently of the level of intensity, making it overly challenging for the model to rank the crops effectively. Moreover, considering only $L\!=\!3$ levels of degradation provides less information to the model during training compared to using five different levels and thus leads to worse results. Then, we observe that directly employing the predicted quality scores in the ranking loss instead of the similarity to the prompts achieves poor performance. We attribute this outcome to an increased discrepancy between CLIP’s training and fine-tuning process. Indeed, while the predicted quality scores originate from two prompts (see Eq. 5), the proposed strategy considers multiple pairs of single images and texts, which we argue is more similar to the technique used for training CLIP [28]. Finally, employing an interpolated positional embedding produces comparable results to its removal. This contrasts with observations from the authors of CLIP-IQA [41], who reported a performance drop when the positional embedding was used. We suppose that in our case the model adapts to the presence or absence of the positional embedding during the fine-tuning process, thus leading to similar outcomes.
Figure 4: gMAD competition results between QualiCLIP and GRepQ [36]. (a): Fixed QualiCLIP at a low- (top) and high-quality (bottom) level, respectively. (b): Fixed GRepQ at a low- (top) and high-quality (bottom) level, respectively.
4.3 Additional Experiments
We evaluate the robustness and the explainability of our model through the gMAD [22] competition and a gradCAM [32] visualization, respectively. We report additional experiments in the supplementary material.
gMAD To assess the robustness of our model we carry out the group maximum differentiation (gMAD) competition [22]. In particular, we compare QualiCLIP against GRepQ using the Waterloo Exploration Database [21] dataset, which comprises 95K synthetically degraded images without MOS annotations. In this evaluation, one model is fixed to function as a defender, and its quality predictions are grouped into two distinct levels. The other model assumes the role of the attacker, tasked with identifying image pairs within each level that exhibit the greatest quality difference. For a model to demonstrate robustness, the selected image pairs should show comparable quality when acting as the defender while exhibiting a notable quality disparity when assuming the role of the attacker. We observe that when we fix QualiCLIP at a low-quality level (Fig. 4(a) top), GRepQ fails to find picture pairs with an obvious quality difference. When considering a high-quality level (Fig. 4(a) bottom), the image pair identified by GRepQ shows a slight quality gap. However, when assuming the role of the attacker (Fig. 4(b)), QualiCLIP successfully exposes the failures of GRepQ, as it pinpoints image pairs displaying a significant quality disparity. Hence, our approach demonstrates greater robustness than GRepQ.
(a) Positive prompt
(b) Negative prompt
Figure 5: gradCAM visualization of the most important regions of the input image for each of the antonym prompts.
gradCAM visualization We evaluate the explainability of our model and CLIP-IQA via a gradCAM [32] visualization. gradCAM is a visualization technique aimed at understanding which regions of an input image are most influential for a model’s decision by studying the gradients of a given layer. We employ gradCAM to produce a heatmap of the regions of the image that activate the most for each of the antonym prompts. We employ “Good photo” and “Bad photo” as the positive and negative prompts, respectively. Following [32], we consider the last convolutional layer of the ResNet50 backbone. Fig. 5(a) shows the result for the positive prompt. We observe that, compared to CLIP-IQA, our model leads to a better alignment with high-quality areas of the image, such as the head of the horse. Similarly, Fig. 5(b) illustrates that QualiCLIP focuses on the most degraded parts of the images when considering the negative prompt, in contrast with CLIP-IQA. This experiment shows that our training strategy forces CLIP to focus on the low-level characteristics of the images. Moreover, the improved alignment between the antonym prompts and the corresponding regions of the images makes QualiCLIP more easily explainable than CLIP-IQA.
@@ 5 Conclusion
In this work, we propose QualiCLIP, a self-supervised opinion-unaware approach aimed at enhancing CLIP’s ability to produce accurate quality-aware image representations. In particular, we design a quality-aware image-text alignment strategy that trains CLIP to rank increasingly synthetically degraded images based on their similarity with antonym prompts, while ensuring consistent representations for images with similar content and comparable quality. The experiments show that QualiCLIP outperforms existing opinion-unaware methods – up to a 20% relative gain – and, in most cases, even supervised approaches in the cross-dataset setting. For this reason, we believe that our method could serve as a strong baseline for evaluating the generalization capabilities of future NR-IQA approaches.
@@ Acknowledgments
    Datasets: we report the details about the datasets employed in the experiments;
@@
      Quantitative Results: we provide the zero-shot results on synthetic datasets and the cross-dataset performance when training the baselines on the CLIVE dataset;
@@
      Ablation Studies: we present the ablation results on the authentic datasets, study the impact on the performance of the backbone of CLIP’s image encoder, and analyze the individual prompt contributions;
@@
      gMAD Competition: we conduct the gMAD competition against CLIP-IQA;
@@
We employ both synthetically and authentically degraded datasets to carry out the experiments. Specifically, we consider four synthetic datasets: LIVE [33], CSIQ [14], TID2013 [27], and KADID-10k [17]. LIVE comprises 779 images degraded with 5 different distortion types at 5 levels of intensity, with 29 reference images as the base. CSIQ originates from 30 reference images, each distorted with 6 distinct degradations at 5 intensity levels, resulting in 866 images. TID2013 and KADID-10k comprise 3000 and 10125 images degraded using 24 and 25 types of distortion across 5 different degrees of intensity, originating from 25 and 81 reference images, respectively. We also rely on five authentic datasets: BID [5], KonIQ-10k [12], CLIVE [7], FLIVE [47], and SPAQ [6]. BID comprises 586 images corrupted with 5 different types of realistic blur distortions. KonIQ-10k contains 10K images sampled from the YFCC100M [38] database. CLIVE consists of 1162 images captured with a wide range of mobile devices. FLIVE is the largest existing dataset for NR-IQA and is composed of about 40K real-world images. SPAQ comprises 11K high-resolution photos taken with several smartphones. Following [6], we resize the SPAQ images so that the shorter side is 512 pixels.
@@
LIVE CSIQ TID2013 KADID-10k Average
Method SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
NIQE [26] 0.908 0.905 0.628 0.719 0.312 0.398 0.379 0.438 0.557 0.615
IL-NIQE [49] 0.896 0.896 0.824 0.860 0.487 0.582 0.539 0.582 0.687 0.730
CONTRIQUE-OU [24] 0.854 0.848 0.695 0.715 0.323 0.360 0.552 0.563 0.606 0.622
Re-IQA-OU [31] 0.803 0.796 0.719 0.727 0.288 0.326 0.518 0.531 0.582 0.595
ARNIQA-OU [1] 0.871 0.863 0.816 0.805 0.464 0.533 0.630 0.635 0.695 0.709
CL-MI [3] 0.748 0.732 0.588 0.589 0.253 0.316 0.506 0.513 0.524 0.538
CLIP-IQA [41] 0.663 0.663 0.723 0.781 0.504 0.600 0.480 0.485 0.593 0.632
GRepQ-OU [36] 0.727 0.717 0.692 0.706 0.402 0.550 0.423 0.471 0.561 0.611
QualiCLIP 0.887 0.880 0.772 0.812 0.626 0.679 0.655 0.660 0.735 0.758
Table S1: Comparison between QualiCLIP and competing opinion-unaware methods on datasets with synthetic distortions. Best and second-best scores are highlighted in bold and underlined, respectively. OU indicates Opinion-Unaware version as explained in Sec. 4.1.
Opinion BID KonIQ-10k FLIVE SPAQ Average
Method Unaware SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
HyperIQA [37] 0.628 0.627 0.750 0.787 0.335 0.483 0.776 0.796 0.622 0.673
TReS [8] 0.634 0.638 0.738 0.766 0.356 0.477 0.865 0.870 0.648 0.688
CONTRIQUE [24] 0.640 0.641 0.734 0.747 0.355 0.465 0.840 0.850 0.642 0.676
Re-IQA [31] 0.611 0.609 0.732 0.753 0.341 0.449 0.823 0.832 0.627 0.661
ARNIQA [1] 0.661 0.672 0.751 0.781 0.384 0.480 0.862 0.872 0.665 0.701
CLIP-IQA+ [41] 0.636 0.640 0.780 0.814 0.369 0.481 0.855 0.861 0.660 0.705
LIQE [51] 0.611 0.597 0.809 0.795 0.290 0.353 0.832 0.846 0.636 0.648
GRepQ [36] 0.510 0.499 0.779 0.793 0.345 0.449 0.839 0.852 0.618 0.648
QualiCLIP 0.646 0.657 0.817 0.841 0.392 0.499 0.848 0.857 0.676 0.714
Table S2: Comparison between QualiCLIP and supervised methods trained on the CLIVE dataset. We report the performance on several datasets with authentic distortions. Best and second-best scores are highlighted in bold and underlined, respectively.
@@
Zero-shot setting We compare the performance of our model on synthetic datasets against existing opinion-unaware approaches in Tab. S1. Note that LIVE and TID2013 were used for validation, thus potentially introducing some bias in our results. Nevertheless, we report them for completeness. Although QualiCLIP is primarily designed for authentically distorted images in real-world scenarios, it achieves state-of-the-art performance on synthetic datasets as well. Indeed, consistent with our observations on authentic datasets in Sec. 4.1, our method significantly outperforms all the baselines on average.
Cross-dataset setting Tab. S2 shows the results for the cross-dataset setting when using the CLIVE [7] dataset for training the baselines. Despite being an opinion-unaware approach, QualiCLIP obtains state-of-the-art results. In particular, our method achieves superior average performance than other CLIP-based methods, namely CLIP-IQA+ [41], LIQE [51] and GRepQ [36]. This outcome aligns with the results reported in Tab. 2 and further confirms the effectiveness of our quality-aware image-text alignment strategy.
@@
$\mathcal{L}_{cons}$ $\mathcal{L}_{pos}$ $\mathcal{L}_{neg}$ BID KonIQ CLIVE FLIVE SPAQ
0.635 0.734 0.656 0.346 0.743
0.636 0.760 0.640 0.337 0.768
0.638 0.801 0.735 0.386 0.810
0.636 0.767 0.658 0.348 0.791
0.641 0.808 0.742 0.389 0.828
0.643 0.805 0.738 0.388 0.821
0.646 0.817 0.756 0.392 0.848
Ablation BID KonIQ CLIVE FLIVE SPAQ
D = 2 0.645 0.812 0.734 0.378 0.835
L = 3 0.643 0.808 0.733 0.367 0.792
quality-based 0.632 0.748 0.704 0.356 0.756
w/ pos. emb. 0.647 0.816 0.742 0.385 0.846
QualiCLIP 0.646 0.817 0.756 0.392 0.848
Table S3: Ablation studies on the loss terms of Eq. 4 (top) and on the training strategy (bottom). Best and second-best SRCC scores are highlighted in bold and underlined, respectively.
Losses and training strategy
In Sec. 4.2 we conduct ablation studies on synthetic datasets, reserving authentic ones exclusively for testing to emulate real-world scenarios. Nevertheless, for completeness, we report the results on authentic datasets for different variants of our approach in Tab. S3. The same considerations made for synthetic datasets apply to authentic ones. Specifically, the top part of Tab. S3 shows that $\mathcal{L}_{neg}$ (Eq. 3) is the most critical component of the training loss and that the combination of the three loss terms $\mathcal{L}_{cons}$ (Eq. 1), $\mathcal{L}_{pos}$ (Eq. 2), and $\mathcal{L}_{neg}$ achieves the best performance. The bottom section of Tab. S3 illustrates that using more sequential degradations, fewer levels of intensity, and relying on quality scores instead of the similarity between images and antonym prompts hinders the performance. Finally, we observe that interpolating the positional embedding instead of removing it yields comparable results.
Backbone of CLIP’s image encoder
Following [41], we evaluate the impact of the backbone architecture of CLIP’s image encoder on performance. Specifically, we examine the ResNet50 and ViT-B/32 backbones. In Tab. S4 we compare the results of QualiCLIP with CLIP-IQA [41], which leverages an off-the-shelf CLIP model. As also observed by Wang et al. [41], relying on the ViT-B/32 significantly hinders the performance over the ResNet50. This outcome stems from the stronger inductive bias of convolutional networks compared to transformers, which are more sensitive to the removal of the positional embedding. Nevertheless, for both backbones, we observe that QualiCLIP significantly outperforms CLIP-IQA.
Method BID KonIQ CLIVE FLIVE SPAQ
RN50 CLIP-IQA 0.634 0.699 0.611 0.287 0.733
QualiCLIP 0.646 0.817 0.756 0.392 0.848
B/32 CLIP-IQA 0.351 0.381 0.376 0.230 0.603
QualiCLIP 0.428 0.625 0.395 0.237 0.784
@@
Analysis of individual prompt contributions
The results of the ablation studies on the training loss terms reported in Sec. 4.2 and Tab. S3 show that $\mathcal{L}_{neg}$ is more crucial than $\mathcal{L}_{pos}$ for the training process. We recall that $\mathcal{L}_{pos}$ and $\mathcal{L}_{neg}$ involve the alignment between the images and the positive and negative prompts, respectively. Therefore, this finding suggests that the negative prompt has a greater influence than the positive one on the quality score computation (as in Eq. 5). To support this hypothesis, we study the individual contribution of the positive and negative prompts in obtaining the final quality scores.
Let $T_{p}$ and $T_{n}$ be the positive and negative prompts that compose the pair of antonym prompts. We perform an experiment where we directly use the similarity between the image and each of the antonym prompts as the quality score. This is possible because both the similarities and the quality scores range between 0 and 1. Tab. S5 shows the results on the authentic datasets. We observe that the similarity between the negative prompt and the image provides significantly more information about its inherent quality compared to using the positive prompt. This result confirms our hypothesis and is consistent with the greater importance of $\mathcal{L}_{neg}$ in our training strategy. Nevertheless, Tab. S5 also indicates that both the positive and negative prompts are crucial for the quality score computation, as the strategy detailed in Eq. 5 achieves the best performance.
We carry out an additional experiment to investigate whether the discrepancy in the contribution between the positive and negative prompts is a result of our training strategy or is inherent to CLIP itself. Specifically, we follow the experimental setting described above to evaluate the individual contributions of the prompts in the quality score computation of CLIP-IQA [41]. We recall that CLIP-IQA computes the quality score with a strategy similar to Eq. 5. Our experiment reveals that $T_{p}$ and $T_{n}$ achieve a SRCC of $0.011$ and $0.578$ on the KonIQ-10k dataset, respectively. This outcome leads us to conclude that the similarity with the negative prompt inherently provides more meaningful information about image quality compared to using the positive prompt. We will investigate this finding more thoroughly in future work.
$T_{p}$ $T_{n}$ BID KonIQ CLIVE FLIVE SPAQ
0.259 0.219 0.060 0.010 0.041
0.550 0.664 0.678 0.311 0.816
0.646 0.817 0.756 0.392 0.848
Table S5: Analysis of the individual prompt contributions in the quality score computation. $T_{p}$ and $T_{n}$ indicate the positive and negative prompts, respectively. Best and second-best SRCC scores are highlighted in bold and underlined, respectively.

### v3

@@ 4.2 Ablation Studies
We conduct ablation studies to analyze the impact of different components of our training strategy, the importance of each loss term, and the contribution of each of the antonym prompts in the quality score computation. For simplicity, we only report the SRCC results on the authentic datasets.
Ablation KonIQ-10k CLIVE FLIVE SPAQ
D = 2 0.815 0.718 0.437 0.834
L = 3 0.808 0.719 0.382 0.788
quality-based 0.745 0.708 0.396 0.756
QualiCLIP 0.817 0.725 0.442 0.841
@@
Table 3: Ablation study on the training strategy. Best and second-best scores are highlighted in bold and underlined, respectively.
@@
Training strategy We evaluate the performance achieved by modified versions of our approach: 1) $D\!=\!2$: we apply two sequential degradations to each crop in Sec. 3.2 instead of just one; 2) $L\!=\!3$: we consider only three levels of degradation in Secs. 3.2 and 3.3 instead of five; 3) we compute the ranking loss using the predicted quality scores – obtained with Eq. 5 – associated to each degraded crop instead of its similarity to the antonym prompts; Tab. 3 shows the results. First, we notice that employing more than one distortion leads to slightly worse performance. We argue that this result stems from the synthetic degradation becoming too severe independently of the level of intensity, making it overly challenging for the model to rank the crops effectively. Moreover, considering only $L\!=\!3$ levels of degradation provides less information to the model during training compared to using five different levels and thus significantly worsens the results. Then, we observe that directly employing the predicted quality scores in the ranking loss instead of the similarity to the prompts achieves poor performance. We attribute this outcome to an increased discrepancy between CLIP’s training and fine-tuning process. Indeed, while the predicted quality scores originate from two prompts (see Eq. 5), the proposed strategy considers multiple pairs of single images and texts, which we argue is more similar to the technique used for training CLIP [32].
$\mathcal{L}_{cons}$ $\mathcal{L}_{pos}$ $\mathcal{L}_{neg}$ KonIQ-10k CLIVE FLIVE SPAQ
0.733 0.668 0.368 0.703
0.764 0.532 0.410 0.769
0.805 0.714 0.420 0.812
0.771 0.550 0.404 0.791
0.800 0.715 0.427 0.820
0.811 0.710 0.434 0.837
0.817 0.725 0.442 0.841
Table 4: Ablation study on the loss terms. Best and second-best scores are highlighted in bold and underlined, respectively.
Loss terms We study the importance of each loss term in Eq. 4 and report the results in Tab. 4. First, we notice that using only $\mathcal{L}_{cons}$ leads to a significant performance decrease, as $\mathcal{L}_{cons}$ does not exploit the information provided by the intrinsic ranking of the increasingly degraded crops. Nevertheless, $\mathcal{L}_{cons}$ consistently yields a positive impact when combined with any of the other loss terms. Then, we observe that, while $\mathcal{L}_{pos}$ and $\mathcal{L}_{neg}$ differ only for the type of prompt they consider, $\mathcal{L}_{neg}$ proves to be significantly more critical for the training process. Nevertheless, Tab. 4 shows that combining the three loss terms achieves the best results, proving that they are all crucial for training CLIP to generate accurate quality-aware image representations.
Individual prompt contributions The results of the ablation studies on the training loss terms show that $\mathcal{L}_{neg}$ is more critical than $\mathcal{L}_{pos}$ for the training process. We recall that $\mathcal{L}_{pos}$ and $\mathcal{L}_{neg}$ involve the alignment between the images and the positive and negative prompts, respectively. This suggests that the similarity between the image and the negative prompt has a greater influence than that with the positive prompt on the quality score computation (as in Eq. 5). To support this hypothesis, we study the individual prompt contributions in obtaining the final quality scores.
We conduct an experiment where we directly use the similarity between the image and each of the antonym prompts as the quality score. This is possible because both the similarities and the quality scores range between 0 and 1. Tab. 5 shows the results. We observe that the similarity between the negative prompt and the image provides significantly more information about its inherent quality than the positive prompt. This result supports our hypothesis and is consistent with the greater importance of $\mathcal{L}_{neg}$ in our training strategy. Nonetheless, Tab. 5 also indicates that both prompts are essential for the quality score computation, as their combination results in the best performance.
We carry out an additional experiment to determine whether the discrepancy in the contributions of the positive and negative prompts arises from our training strategy or is inherent to CLIP itself. Specifically, we follow the experimental setting described above to evaluate the individual contributions of the prompts in the quality score computation of CLIP-IQA [47]. We recall that CLIP-IQA employs an off-the-shelf CLIP model and computes the final quality scores using a strategy similar to Eq. 5. Our experiment reveals that using $T_{p}$ and $T_{n}$ individually results in a SRCC of $0.010$ and $0.571$, respectively, on the KonIQ-10k dataset. This outcome leads us to conclude that the similarity with the negative prompt inherently provides more meaningful information about image quality compared to using the positive prompt. We will investigate this finding more thoroughly in future work.
$T_{p}$ $T_{n}$ KonIQ-10k CLIVE FLIVE SPAQ
0.230 0.063 0.064 0.059
0.661 0.686 0.332 0.818
0.817 0.725 0.442 0.841
Table 5: Analysis of the individual prompt contributions in the quality score computation. Best and second-best scores are highlighted in bold and underlined, respectively.
@@ 5 Conclusion
In this work, we propose QualiCLIP, a self-supervised opinion-unaware approach that enhances CLIP’s ability to produce accurate quality-aware image representations. In particular, we design a quality-aware image-text alignment strategy that trains CLIP to rank increasingly synthetically degraded images based on their similarity with antonym prompts, while ensuring consistent representations for images with similar content and comparable quality. Compared to existing opinion-unaware methods, QualiCLIP shows significant performance improvements across several datasets. Moreover, it is the only opinion-unaware approach that, in most cases, outperforms opinion-aware methods in cross-dataset experiments. Thus, we believe that QualiCLIP could serve as a strong baseline for evaluating the generalization capabilities of future NR-IQA approaches.
@@ Acknowledgments
    Datasets: we report the details of the datasets employed in the experiments;
@@
      Quantitative Results: we provide the zero-shot results on synthetic datasets and the cross-dataset performance using the CLIVE and PIPAL datasets for training the baselines;
@@
      Ablation Studies: we analyze the impact on the performance of the backbone of CLIP’s image encoder;
@@
      gMAD Competition: we conduct the gMAD competition against GRepQ and CLIP-IQA;
@@
      Supervised QualiCLIP: we extend our approach to leverage human annotations;
    6. S2.6.
@@
To carry out the experiments, we employ several types of datasets, namely authentic, image restoration, AIGC, and synthetic datasets. We rely on four authentic datasets: KonIQ-10k [11], CLIVE [6], FLIVE [53], and SPAQ [5]. KonIQ-10k contains 10K images sampled from the YFCC100M [43] database. CLIVE consists of 1162 images captured with a wide range of mobile devices. FLIVE is the largest existing dataset for NR-IQA and is composed of about 40K real-world images. SPAQ comprises 11K high-resolution photos taken with several smartphones. Following [5], we resize the SPAQ images so that the shorter side is 512 pixels. We use two image restoration datasets: CVIU [22] and PIPAL [12]. CVIU stems from 30 reference images distorted with 9 super-resolution methods, resulting in 1620 images. PIPAL comprises 23200 images degraded with 40 distortion types, including GAN-based super-resolution methods, and originates from 250 reference images. We employ two AIGC datasets: AGIQA-1K [59] and AGIQA-3K [15]. AGIQA-1K includes 1080 images generated with 2 diffusion models. AGIQA-3K consists of 2982 images generated via 6 generative models, including auto-regressive and diffusion ones. We consider four synthetic datasets: LIVE [37], CSIQ [14], TID2013 [31], and KADID-10k [18]. LIVE comprises 779 images degraded with 5 different distortion types at 5 levels of intensity, with 29 reference images as the base. CSIQ originates from 30 reference images, each distorted with 6 distinct degradations at 5 intensity levels, resulting in 866 images. TID2013 and KADID-10k comprise 3000 and 10125 images degraded using 24 and 25 types of distortion across 5 different degrees of intensity, originating from 25 and 81 reference images, respectively.
@@
LIVE CSIQ TID2013 KADID-10k
Method OU SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
CONTRIQUE [26] 0.962 0.967 0.946 0.960 0.821 0.843 0.927 0.928
Re-IQA [35] 0.954 0.956 0.946 0.954 0.873 0.895 0.867 0.865
ARNIQA [1] 0.965 0.970 0.961 0.969 0.874 0.896 0.905 0.908
CLIP-IQA+ [47] 0.950 0.946 0.862 0.882 0.804 0.834 0.818 0.822
GRepQ [41] 0.926 0.931 0.844 0.871 0.668 0.681 0.808 0.797
NIQE [29] 0.903 0.906 0.642 0.727 0.308 0.419 0.380 0.442
IL-NIQE [55] 0.878 0.884 0.848 0.883 0.504 0.636 0.557 0.599
CL-MI [3] 0.749 0.731 0.619 0.615 0.249 0.317 0.518 0.529
MDFS [30] 0.926 0.930 0.797 0.829 0.554 0.657 0.608 0.634
CONTRIQUE-OU [26] 0.841 0.840 0.684 0.700 0.321 0.344 0.552 0.564
Re-IQA-OU [35] 0.781 0.779 0.713 0.725 0.284 0.339 0.516 0.534
ARNIQA-OU [1] 0.853 0.852 0.832 0.810 0.467 0.527 0.632 0.638
CLIP-IQA [47] 0.626 0.651 0.748 0.805 0.525 0.632 0.465 0.473
GRepQ-OU [41] 0.714 0.698 0.740 0.742 0.423 0.565 0.416 0.463
QualiCLIP 0.898 0.885 0.804 0.842 0.651 0.732 0.654 0.664
[…]
Authentic Datasets Image Restoration Datasets AIGC Datasets
KonIQ-10k FLIVE SPAQ CVIU PIPAL AGIQA-1K AGIQA-3K
Method OU SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
HyperIQA [42] 0.734 0.773 0.328 0.483 0.771 0.796 0.632 0.642 0.361 0.384 0.183 0.351 0.594 0.686
DBCNN [57] 0.703 0.755 0.299 0.435 0.806 0.816 0.728 0.748 0.332 0.341 0.445 0.693 0.448 0.558
TReS [7] 0.727 0.756 0.359 0.494 0.858 0.863 0.713 0.739 0.397 0.453 0.364 0.643 0.549 0.666
LIQE [58] 0.810 0.787 0.277 0.271 0.830 0.840 0.704 0.717 0.446 0.445 0.417 0.609 0.653 0.732
QCN [39] 0.762 0.835 0.428 0.537 0.860 0.867 0.765 0.770 0.393 0.413 0.337 0.677 0.584 0.729
CONTRIQUE [26] 0.736 0.749 0.399 0.527 0.834 0.842 0.583 0.601 0.310 0.370 0.256 0.393 0.503 0.597
Re-IQA [35] 0.600 0.618 0.252 0.341 0.820 0.829 0.576 0.685 0.360 0.386 0.330 0.548 0.431 0.516
ARNIQA [1] 0.745 0.775 0.434 0.526 0.861 0.871 0.610 0.643 0.369 0.382 0.320 0.629 0.526 0.649
CLIP-IQA+ [47] 0.769 0.808 0.410 0.547 0.846 0.854 0.837 0.851 0.369 0.385 0.395 0.584 0.637 0.712
GRepQ [41] 0.742 0.750 0.391 0.464 0.855 0.865 0.701 0.717 0.368 0.385 0.346 0.589 0.616 0.730
QualiCLIP 0.817 0.838 0.442 0.556 0.841 0.851 0.840 0.852 0.412 0.422 0.736 0.805 0.667 0.735
Table S2: Quantitative results of the cross-dataset evaluation setting. We employ the CLIVE [6] dataset to train the supervised methods. OU stands for Opinion-Unaware. Best and second-best scores are highlighted in bold and underlined, respectively.
Authentic Datasets IR Dataset AIGC Datasets
KonIQ-10k CLIVE FLIVE SPAQ CVIU AGIQA-1K AGIQA-3K
Method OU SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
HyperIQA [42] 0.570 0.564 0.560 0.642 0.246 0.310 0.752 0.762 0.821 0.835 0.493 0.486 0.497 0.535
DBCNN [57] 0.718 0.744 0.570 0.638 0.267 0.378 0.779 0.795 0.803 0.814 0.427 0.651 0.516 0.601
LIQE [58] 0.681 0.696 0.657 0.691 0.138 0.178 0.692 0.690 0.797 0.810 0.413 0.465 0.560 0.621
QCN [39] 0.584 0.663 0.463 0.559 0.292 0.408 0.785 0.797 0.766 0.767 0.484 0.679 0.563 0.650
CONTRIQUE [26] 0.445 0.460 0.342 0.408 0.228 0.308 0.596 0.600 0.719 0.723 0.311 0.535 0.344 0.392
Re-IQA [35] 0.351 0.376 0.297 0.358 0.206 0.273 0.633 0.641 0.670 0.692 0.291 0.410 0.317 0.375
ARNIQA [1] 0.246 0.325 0.389 0.486 0.224 0.252 0.614 0.644 0.757 0.768 0.173 0.564 0.557 0.629
CLIP-IQA+ [47] 0.568 0.586 0.551 0.597 0.231 0.313 0.758 0.766 0.748 0.764 0.442 0.560 0.679 0.720
GRepQ [41] 0.679 0.682 0.536 0.583 0.285 0.345 0.754 0.765 0.792 0.820 0.174 0.591 0.573 0.679
QualiCLIP 0.817 0.838 0.725 0.802 0.442 0.556 0.841 0.851 0.840 0.852 0.736 0.805 0.667 0.735
Table S3: Quantitative results of the cross-dataset evaluation setting. We employ the PIPAL [12] dataset to train the supervised methods. OU stands for Opinion-Unaware. IR signifies Image Restoration. Best and second-best scores are highlighted in bold and underlined, respectively.
@@
Zero-shot setting In Tab. S1, we compare the performance of our model with existing opinion-unaware approaches on synthetic datasets. We observe that our method achieves competitive performance, obtaining the most consistent results among the considered approaches. However, as also observed in Sec. 4.1, supervised opinion-aware methods achieve better performance than opinion-unaware ones when a training set is available.
Cross-dataset setting Tabs. S2 and S3 present the results of the cross-dataset setting when employing CLIVE [6] and PIPAL [12], respectively, to train the supervised baselines. Note that we do not report the performance of TReS [7] with PIPAL as the training dataset as there is no public pre-trained model available. The results show that QualiCLIP achieves excellent performance regardless of the dataset used to train the baselines. Moreover, we observe that the opinion-aware approaches generally perform worse when trained on PIPAL compared to CLIVE. We attribute this outcome to the nature of the degradation types included in PIPAL, which are different from those in the other testing datasets. This result highlights the sensitivity of the supervised opinion-aware methods to the training data, and further confirms their limited generalization capabilities.
@@
Backbone of CLIP’s image encoder Following [47], we evaluate the impact of the backbone architecture of CLIP’s image encoder on performance. Specifically, we examine the ResNet50 and ViT-B/32 backbones. In Tab. S4 we compare the results of QualiCLIP with CLIP-IQA [47], which leverages an off-the-shelf CLIP model. As also observed by Wang et al. [47], relying on the ViT-B/32 significantly hinders the performance over the ResNet50. This outcome stems from the stronger inductive bias of convolutional networks compared to transformers, which are more sensitive to the removal of the positional embedding. Nevertheless, for both backbones, we observe that QualiCLIP outperforms CLIP-IQA.
Method KonIQ-10k CLIVE FLIVE SPAQ
RN50 CLIP-IQA 0.695 0.656 0.344 0.733
QualiCLIP 0.817 0.725 0.442 0.841
B/32 CLIP-IQA 0.390 0.328 0.211 0.608
QualiCLIP 0.626 0.334 0.217 0.782